8:30: start of calibration unit activity, spots are already in LGSW

HEXA X Y Z Rx Ry
  -5.02 -7.27 4.38 -365 135

We close the loop without disturb, using TCoG48, and we acquire the new position of the Hexapod once the 3 LGS are centered on the FoV minimizing the tilt on the 3 subaps patterns:

DoF X Y Z Rx Ry
HEXA -5.02 -7.27 4.33 -363 139
TS offload 0.0 0.0 -0.05 1.64 4.10

We switch on the on-axis surce, lamp steps @ 55

TN Notes
20141127_165642  
20141127_165826

We load the pupil target 20141126_183730 and we close the pupil loop.

TN Notes
20141127_170653  
20141127_170734

We set the jitter loop gain to 0.01 and we close

Pupil PI X PI Y
Blue -0.83 2.30
Yellow -1.20 -0.85
Red 2.40 -1.80

TN Notes
20141127_173430 Jitter is drifting because the TT loop is also closed on LGSW
20141127_173613

PBL with DSM: when we send the closeAO command the shell RIPs because of housekeeping errors... Probably related to ARGOS Arbitrator connection through ICE interface.

TN Notes
20141127_180117  

We build a mixed FLAO - ARGOS reconstructor.

Combined reconstructor with FLAO BIN2

COMB REC TAG ARGOS REC FLAO REC COMB REC
TCoG48 + bin2 20141126_215200 20130916_221257 20141127_180600
CoG + bin 2 20141126_210900 20141127_180700
TCoG48 only 20141126_215200 - 20141127_185500
CoG only 20141126_210900 - 20141127_185600
FLAO - 20130916_221257 20141127_200600

Combined RECs test

We close the loop with TCoG48 bin2 REC, PCL, JOL, FLAO is running at 200Hz

TN Notes
20141127_190210 G=0.5,0.4,0.3
20141127_190302
20141127_190341

We activate the TS loop, REC for TS is Rec_20130916_221257_argos.

TN Notes
20141127_192528 G=0.5,0.4,0.3, stage Z = 36mm
20141127_192921 stage Z = 33mm, modulation = 5V
20141127_193142 modulation = 4V

We try to stabilize the PSF on CCD47, we have to rise the TT gain up to 20!

TN Notes
20141127_195627 G=20.0,0.4,0.3

PBL with DSM: having restarted the Arbitrator and closing again the AO loop the shell RIPs again for housekeeping error...

PBL with BCU it stacks when framerate is > 500Hz...

Pupil PI X PI Y
Blue -1.61 0.41
Yellow -0.63 -2.53
Red -1.22 -2.20

Now we move the FLAO Z stage @ 42mm and we close the AO loop just on FLAO. Since the FLAO slopes are still passing through the ARGOS BCU also this loop is running at 500Hz.

TN Notes
20141127_210846 POL
20141127_210906

SLOPE_OFFSET: 20141127_210435: taken w/o focus optimization on CCD47, PCL

SLOPE_OFFSET: 20141127_210509, 20141127_210544: taken after focus optimization on CCD47, Z stage @ 33mm, PCL

SLOPE_OFFSET: 20141127_210730, 20141127_210757: taken after focus optimization on CCD47, Z stage @ 33mm, POL

PUPIL TARGET: 20141127_211425, 20141127_211808

We close the AO loop on FLAO + ARGOS * TS (20 modes)

TN Notes
20141127_214629  
20141127_214710
20141127_214933
20141127_215025
20141127_215311

PUPIL TARGET: 20141127_215903

Now we enable the jitter loop

TN Notes
20141127_220420 JCL, G=0.01
20141127_220457
20141127_220628 JCL, G=0.1
20141127_220707

Focus optimization on LUCI

We scan the focus moving the Z stage on FLAO board, running the TS script.

Stage Z TN Notes
25 20141127_221023  
26 20141127_221119  
27 20141127_221302  
28 20141127_221406  
29 20141127_221503  
30 20141127_221605  
31 20141127_221707  
32 20141127_221759  
33 20141127_221859  
34 20141127_221948  
35 20141127_222040  

Combined reconstructor with FLAO BIN4

COMB REC TAG ARGOS REC FLAO REC COMB REC
TCoG48 + bin4 20141126_215200 20130917_214654 20141127_222222
CoG + bin4 20141126_210900 20141127_222400

We close the TT loop on FLAO bin4 @ 250Hz with TS and ARGOS correcting 150 modes. Stage Z is at 28.5mm.

TN Notes
20141127_230834 G=6,0.3,0.2
20141127_230926
20141127_231020
20141127_231410 Disturb = 1.6", G=6,0.4,0.4, POL, Gj=0.1
20141127_231552 G=6,0.4,0.4, PCL, Gj=0.1  
20141127_231712 G=6.0.4,0.4, Gj=0.2
20141127_231819 G=6.0.4,0.4, Gj=0.3
20141127_232029 G=6.0.4,0.4, Gj=0.4
20141127_232118 G=6.0.4,0.4, Gj=0.5

We ramp up the jitter gain up to 0.7 and then we brake everything...

We restart with PCL and JCL, AO loop is open disturb = 1.6":

TN Notes
20141127_232336 OL ,Gj=0.6
20141127_233311 OL
20141127_233352 OL ,Gj=0.5
20141127_233431 OL ,Gj=0.2

Rotation of PatCams interaction matrices

Since the 3 Patrol Cameras have been dismounted after the August LGSW removal we have measured the new orientation:

PatCam Mar Nov Delta
Blue 0 184 +184
Yellow -150 -158 -8
Red 160 170 +10

So we restored the March reconstructors for LM1-PatCam and periscope-PatCam algorithms and we applied a rotation:

Algo Mar Nov (cwise) Nov (ccwise)
LM1 20140323_033220 20141128_021900 20141128_ 025000
Per 20140322_072532 20141128_022000 20141128_025100
-- MarcoBonaglia - 09 Dec 2014
Topic revision: r4 - 15 Dec 2014, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback